A committee of 5 people is to be formed from 11 civil employees and 8 small business people. Calculate the following :

a) The probability that none of the people selected are small business people
b) The probability at least 2 people are small business people
c) The expected number of small business people on the committee
1
Expert's answer
2020-03-17T14:46:30-0400

Let "X=" the number of small business people on the committee.

There are 19C5 ways to select 5 people (committee members) out of a total of 19 people (11 civil employees and 8 small business people).


"\binom{19}{5}={19! \over 5!(19-5)!}={19(18)(17)(16)(15) \over 1(2)(3)(4)(5)}=11628"



a) The probability that none of the people selected are small business people  


"P(X=0)={ \dbinom{11}{5-0}\dbinom{8}{0}\over \dbinom{19}{8}}=\dfrac{\dfrac{11!}{5!(11-5)!}\cdot\dfrac{8!}{0!(8-0)}}{11628}=""={462\cdot1 \over 11628}={77 \over 1938}\approx0.0397"

b) The probability at least 2 people are small business people  


"P(X\geq2)=1-P(X=0)-P(X=1)=""=1-{ \dbinom{11}{5-0}\dbinom{8}{0}\over \dbinom{19}{8}}-{ \dbinom{11}{5-1}\dbinom{8}{1}\over \dbinom{19}{8}}=""=1-\dfrac{\dfrac{11!}{5!(11-5)!}\cdot\dfrac{8!}{0!(8-0)}}{11628}-\dfrac{\dfrac{11!}{4!(11-4)!}\cdot\dfrac{8!}{1!(8-1)}}{11628}="

"=1-{462\cdot1 \over 11628}-{330\cdot8 \over 11628}={1421 \over 1938}\approx0.7332"

c) The expected number of small business people on the committee.


"P(X=0)={77 \over 1938}"

"P(X=1)={440 \over 1938}"

"P(X=2)={ \dbinom{11}{5-2}\dbinom{8}{2}\over \dbinom{19}{8}}=\dfrac{\dfrac{11!}{3!(11-3)!}\cdot\dfrac{8!}{2!(8-2)}}{11628}=""={165\cdot28 \over 11628}={770 \over 1938}"

"P(X=3)={ \dbinom{11}{5-3}\dbinom{8}{3}\over \dbinom{19}{8}}=\dfrac{\dfrac{11!}{2!(11-2)!}\cdot\dfrac{8!}{3!(8-3)}}{11628}=""={55\cdot56 \over 11628}={770 \over 2907}"

"P(X=4)={ \dbinom{11}{5-4}\dbinom{8}{4}\over \dbinom{19}{8}}=\dfrac{\dfrac{11!}{1!(11-1)!}\cdot\dfrac{8!}{4!(8-4)}}{11628}=""={11\cdot70 \over 11628}={385 \over 5814}"

"P(X=5)={ \dbinom{11}{5-5}\dbinom{8}{5}\over \dbinom{19}{8}}=\dfrac{\dfrac{11!}{0!(11-0)!}\cdot\dfrac{8!}{5!(8-5)}}{11628}=""={1\cdot56 \over 11628}={14\over 2907}"

"E(X)=\displaystyle\sum_{i=0}^5x_iP(X=x_i)=0\cdot{77 \over 1938}+1\cdot{440 \over 1938}+"

"+2\cdot{770 \over 1938}+3\cdot{770 \over 2907}+4\cdot{385 \over 5814}+5\cdot{14 \over 2907}={40 \over 19}"

"E(X)={40 \over 19}\approx2.1053"